OpenStudy (anonymous):

A tank in the shape of an inverted right circular cone has height 10 meters and radius 18 meters. It is filled with 5 meters of hot chocolate. Find the work required to empty the tank by pumping the hot chocolate over the top of the tank. Note: the density of hot chocolate is 1490kg/m^3

sam (.sam.):

determine the centroid of inverted circular cone, infinite small of mass, dM = (rho)πr² dy (1) use triangular proportion, r/y = R/H r = Ry/H (2) substitute (2) into (1), dM = (rho)π(Ry/H)² dy static moment with respect to x -axis is Sx = ∫ y dM Y Sx = ∫ y ρπ(Ry/H)² dy 0 Y Sx = ¼ y⁴ ρπ(R/H)² | 0 Sx = ¼ Y⁴ ρπ(R/H)² total mass is, Y M = ∫ ρπ(Ry/H)² dy 0 Y M = ⅓ y³ ρπ(R/H)² | 0 M = ⅓ Y³ ρπ(R/H)² height of inverted right circular cone centroid is Y(o) = Sx/M = ¼ Y⁴ ρπ(R/H)²/{⅓ Y³ ρπ(R/H)²} Y(o)= ¾ Y W=mgh W=(ρ_chocolate)(V_chocolate) g h W=(ρ_chocolate)(⅓ πR² H) (g) Y(o) W=(1540)(⅓ π * (4 * 17/5)² * 4)(9.81)(¾ * 4) W = 8778441.47 J

OpenStudy (anonymous):

W = F * d , integral [ 5, 10 ] = (weight * upper volume ) * displacement = ( 1490 * π r² * dy ) * ( 10 - y ) = [ 1490 * π ( 9y/5)² * dy ] * ( 10 - y ) = 1490 π ∫ ( 9y/5)² ( 10 -y) dy [ 5, 10] = 8,689,056 J